Jump to main content.


Drake Chemical

EPA Daily Situation Reports for March 1999

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Wednesday, March 31, 1999

FOR DATES: From 7:00 a.m. Tuesday, March 30, 1999 to 7:00 a.m. Wednesday, March 31, 1999.

WEATHER: Sunny and clear. The temperature range was 29oF to 62oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 20.4 hours and 665.4 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is primarily attributed to the stoppage of feed in order to repair a gap between the floor plates of the Dry Ash Drag (DAD). The TDF processed contaminated soil at an average rate of 32.6 tons per hour while on-line. The contractor is performing related site activities such as feed preparation, debris separation and equipment demobilization. The second and third run of quarterly stack testing for metals and particulate will be performed on this date.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, weekly NPDES water sampling, wastewater management, feed preparation, debris separation, equipment decontamination, dust control, Concover placement, ash handling and ash backfilling. Excavation activities were performed in the exclusion zone adjacent to the CZ113 sheet piling. The contractor hand picked rock from the screeeners for backfilling on the 2-foot rock bottom. Continued with the real time perimeter air monitoring (PAM). The first run of stack testing for metals and particulate was performed.

Received bottom ash analytical reports (bot-ash-03-26-99-391) for the contents of bin 1, (bot-ash-03-22-99-387 and bot-ash-03-24-99-389) for the contents of bin 3 and (bot-ash-03-23-99-388) for the contents of bin 4. All of the analytical test results received on this date indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Tony Eicher (FOCUS), Sandra Downs and Larry Newcomer (PADEP), Kurt Davis (TAG). A group of Officers from the Newcumberland Army Depot toured the site.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Tuesday, March 30, 1999

FOR DATES: From 7:00 a.m. Monday, March 29, 1999 to 7:00 a.m. Tuesday, March 30, 1999.

WEATHER: Sunny and clear. The temperature range was 29oF to 62oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.6 hours and 699.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 29.7 tons per hour while on-line. The contractor is performing related site activities such as feed preparation, debris separation and equipment demobilization. The final quarterly stack testing event for metals and particulate is scheduled to start during the afternoon of Tuesday, March 30, 1999.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, equipment decontamination, dust control, excavation surveys, ash handling and ash backfilling. The contractor hand picked rock from the screeeners for backfilling on the 2-foot rock bottom. Contaminated soils continued to be removed from above the HDPE liner inside of the feed preparation building. Continued with the real time perimeter air monitoring (PAM).

Received bottom ash analytical report (bot-ash-03-25-99-390) for the contents of bin 2. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Monday, March 29, 1999

FOR DATES: From 7:00 a.m. Sunday, March 28, 1999 to 7:00 a.m. Monday, March 29, 1999.

WEATHER: Sunny and clear. The temperature range was 27oF to 60oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 733.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 30.8 tons per hour while on-line. The contractor is performing related site activities such as feed preparation and equipment demobilization.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, equipment decontamination, dust control, ash handling and ash backfilling. Contaminated soils continued to be removed from above the HDPE liner inside of the feed preparation building. Continued with the real time perimeter air monitoring (PAM).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. Stack testing for metals and particulate will be performed on March 30th and 31st.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None

PREPARED BY: Danny Zion for William Werntges (USACE)


Sunday, March 28, 1999

FOR DATES: From 7:00 a.m. Saturday, March 27, 1999 to 7:00 a.m. Sunday, March 28, 1999.

WEATHER: Sunny and clear. The temperature range was 25oF to 58oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 740.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 31 tons per hour while on-line. The contractor is performing related site activities such as feed preparation and equipment demobilization.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, equipment decontamination, dust control, ash handling and ash backfilling. Contaminated soils continued to be removed from above the HDPE liner inside of the feed preparation building. Continued with the real time perimeter air monitoring (PAM) program and recovered the OSHA method 93 24-hour air-sampling event for BNA at the on-site and off-site air monitoring stations.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. Stack testing for metals and particulate will be performed on March 30th and 31st.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for William Werntges (USACE)


Saturday, March 27, 1999

FOR DATES: From 7:00 a.m. Friday, March 26, 1999 to 7:00 a.m. Saturday, March 27, 1999.

WEATHER: Sunny and clear. The temperature range was 23oF to 50oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 747.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 31.3 tons per hour while on-line. The contractor is performing related site activities such as feed preparation and equipment demobilization.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, equipment decontamination, dust control, lime mixing, decontaminated rock backfilling, ash handling and ash backfilling. Contaminated soils continued to be removed from above the HDPE liner inside of the feed preparation building. Continued with the real time perimeter air monitoring (PAM) program and began the OSHA method 93 24-hour air sampling event for BNA at the on-site and off-site air monitoring stations.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. Stack testing for metals and particulate will be performed on March 30th and 31st .

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Tour group from Penn Technical College.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Friday, March 26, 1999

FOR DATES: From 7:00 a.m. Thursday, March 25, 1999 to 7:00 a.m. Friday, March 26, 1999.

WEATHER: Sunny and clear. Temperature range was 26oF to 46oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 22.4 hours and 660.4 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is primarily attributed to the stoppage of feed in order to remove a piece of metal debris from the Dry Ash Drag (DAD) conveyance system. The TDF processed contaminated soil at an average rate of 29.5 tons per hour while on-line. The contractor is performing related site activities such as feed preparation and debris management.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, equipment decontamination, dust control, lime mixing, shot rock backfilling, ash handling and ash backfilling. Continued to excavate soil inside of the feed preparation building and clean the HDPE liner as part of the project close out plan. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical reports (bot-ash-03-07-99-381A and bot-ash-03-07-99-381B) for the contents of bin 7. Bottom ash analytical reports indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Received fly ash analytical report (fly-ash-03-08-99-076) for the contents of bin D. Fly ash analytical test results indicated that the fly ash materials met the project specifications for restricted backfilling on-site (i.e. above elevation 553) due to the concentrations of cadmium, chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. The final quarterly particulate/metals stack testing event has be scheduled for March 30th and 31st .

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Kurt Davis (TAG).

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Thursday, March 25, 1999

FOR DATES: From 7:00 a.m. Wednesday, March 24, 1999 to 7:00 a.m. Thursday, March 25, 1999.

WEATHER: Partly cloudy. Temperature range was 31oF to 55oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 21.5 hours and 622.9 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is primarily attributed to the stoppage of feed in order to perform maintenance to the rotary kiln's rotational speed sensor. The TDF processed contaminated soil at an average rate of 30 tons per hour while on-line. The contractor is performing related site activities such as feed preparation and debris management.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, equipment decontamination, shot rock backfilling, decontaminated rock backfilling, ash handling and ash backfilling. The utility corridor was repaired and re-graded. Continued to excavate soil inside of the feed preparation building and clean the HDPE liner as part of the project close out plan. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. The final quarterly particulate/metals stack testing event has be scheduled for March 30th and 31st .

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Wednesday, March 24, 1999

FOR DATES: From 7:00 a.m. Tuesday, March 23, 1999 to 7:00 a.m. Wednesday, March 24, 1999.

WEATHER: Partly cloudy. Temperature range was 31oF to 48oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 679.3 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 28.4 tons per hour while on-line. The contractor is performing related site activities such as feed preparation and debris management.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, equipment decontamination, shot rock backfilling, ash handling and ash backfilling. The utility corridor was repaired and re-graded. Excavated the soil inside of the feed preparation building as part of the project close out plan. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. The final quarterly particulate/metals stack testing event has be scheduled for March 30th and 31st .

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Larry Newcomer and Mike Welch (PADEP).

PREPARED BY: Steve Gibson for William Werntges (USACE)


Tuesday, March 23, 1999

No report available.


Monday, March 22, 1999

FOR DATES: From 7:00 a.m. Sunday, March 21, 1999 to 7:00 a.m. Monday, March 22, 1999.

WEATHER: Rain /snow with about 1 inch of snow accumulation. Temperature range was 31oF to 39oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently off-line to allow for the re-installation of the trunion bearing. OHM continues to work 24 hours a day, seven days a week operating and maintaining the TDF. The contractor is performing related site activities such as excavation, debris management and bottom ash backfilling.

II. ACTIONS TAKEN: The contractor continued TDF maintenance and Waste Water Treatment Plant (WWTP) operations. The contractor performed an alignment (tracking) of the kiln while it was cold. Both kiln burners were then re-lit and the re-heating process began in the early evening hours. As part of the project closeout activities the contractor continued repositioning the material handling equipment, exposing the HDPE liner, and stockpiling the removed soil inside of the feed preparation building. Waste feed to the TDF is expected to resume later today. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. The final quarterly particulate/metals stack testing event has be scheduled for March 30th and 31st .

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for Chad Thompson (USACE)


Sunday, March 21, 1999

FOR DATES: From 7:00 a.m. Saturday, March 20, 1999 to 7:00 a.m. Sunday, March 21, 1999.

WEATHER: Cloudy and windy. The temperature range was 27oF to 51oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently off-line to allow for the repair of a damaged trunion bearing shaft. OHM continues to work 24 hours a day, seven days a week operating and maintaining the TDF. The contractor is performing related site activities such as excavation, debris management and bottom ash backfilling.

II. ACTIONS TAKEN: The contractor continued TDF maintenance, including preventative maintenance to the Induced Draft fan and servicing the TDF screw conveyors. In addition, the damaged trunion bearing shaft has been repaired and installed. The contractor also performed Waste Water Treatment Plant (WWTP) operations. Excavation activities were continued inside of the feed preparation building. No bottom ash or fly ash analytical reports were received during this reports coverage. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. Stack testing has been rescheduled for March 30th and 31st.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Richard Lupold for William Werntges (USACE)


Saturday, March 20, 1999

FOR DATES: From 7:00 a.m. Friday, March 19, 1999 to 7:00 a.m. Saturday, March 20, 1999.

WEATHER: Cloudy and windy. The temperature range was 30oF to 44oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently off-line to allow for the repair of a damaged trunion bearing. OHM continues to work 24 hours a day, seven days a week operating and maintaining the TDF. The contractor is performing related site activities such as excavation, debris management and bottom ash backfilling.

II. ACTIONS TAKEN: The contractor continued TDF maintenance, including preventative maintenance to the Induced Draft fan and servicing the TDF screw conveyors. The contractor also performed Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris handling, wood debris management, equipment decontamination, lime mixing, shot rock backfilling, ash backfilling and ash handling. Excavation activities were continued inside of the feed preparation building. As a part of the project close out plan, the power screeners inside of the feed preparation building were temporarily moved and the soil stockpiled beneath them was moved to the exclusion zone. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical reports (bot-ash-03-11-99-383A and bot-ash-03-11-99-383B) for the contents of bin 5, and report (bot-ash-03-14-99-384) for the contents of bin 4. The analytical reports indicate that the thermally treated soil meets the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). These soils will be backfilled above elevation 553.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. Stack testing has been rescheduled for March 30th and 31st.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Friday, March 19, 1999

FOR DATES: From 7:00 a.m. Thursday, March 18, 1999 to 7:00 a.m. Friday, March 19, 1999.

WEATHER: Partly cloudy and windy. The temperature range was 37oF to 56oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently off-line to allow for the repair of a damaged trunion bearing. OHM continues to work 24 hours a day, seven days a week operating and maintaining the TDF. The contractor is performing related site activities such as excavation, debris management and shot rock backfilling.

II. ACTIONS TAKEN: The contractor continued TDF maintenance, including removal and repair of the damaged trunion bearing and the cleaning of the steam scrubber duct work. Maintenance to the quench tower was completed and the Thermal Relief Valve (TRV) was closed at 0024 hours on Friday, March 19, 1999. The repaired trunion bearing is to be returned to the site on Saturday, March 20, 1999, feed is expected to resume early next week.

The contractor also performed Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris handling, wood debris management, equipment decontamination, lime mixing, shot rock backfilling, decontaminated rock backfilling, ash backfilling and ash handling. Excavation activities were continued directly south of the WWTP. Three roll-offs containing PPE and miscellaneous metal were shipped off-site as non-hazardous waste to a hazardous waste landfill. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include bringing the TDF back on-line, operating the TDF, excavation, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. Stack testing for metals and particulate is scheduled for the week of March 29th.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Thursday, March 18, 1999

FOR DATES: From 7:00 a.m. Wednesday, March 17, 1999 to 7:00 a.m. Thursday, March 18, 1999.

WEATHER: Partly cloudy to overcast. The temperature range was 35oF to 66oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently off-line to allow for the repair of a damaged trunion bearing. OHM continues to work 24 hours a day, seven days a week operating and maintaining the TDF. The contractor is performing related site activities such as excavation, debris management and shot rock backfilling.

II. ACTIONS TAKEN: The contractor continued TDF maintenance, including repair operations on the damaged trunion bearing and maintenance to the evaporative cooler. Also, in order to perform maintenance to the Secondary Combustion Chamber (SCC), OHM initiated a planned Thermal Relief Valve (TRV) opening at 2010 hours on Wednesday, March 17, 1999. Feed to the incinerator had been off-line for over 24-hours prior and the proper local authorities were notified of the opening.

The contractor also performed Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris handling, wood debris management, equipment decontamination, lime mixing, shot rock backfilling, decontaminated rock backfilling, excavation survey, ash backfilling and ash handling. Excavation activities were continued directly south of the WWTP. Two roll-offs containing PPE and miscellaneous metal were shipped off-site as non-hazardous waste to a hazardous waste landfill. Continued with the real time perimeter air monitoring (PAM) program. The fourth stack testing event for metals and particulate was canceled because of the damaged trunion bearing and has been re-sheduled for March 30th and 31st.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Tony Eicher (FOCUS), Kurt Davis (AIR).

PREPARED BY: Steve Gibson for William Werntges (USACE)


Wednesday, March 17, 1999

FOR DATES: From 7:00 a.m. Tuesday, March 16, 1999 to 7:00 a.m. Wednesday, March 17, 1999.

WEATHER: Sunny. The temperature range was 29oF to 49oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the and maintaining the Thermal Destruction Facility (TDF). Hazardous waste feed to the TDF has been stopped to allow the contractor to replace one of the rotary kiln's trunion bearings. The TDF was on-line for 4.4 hours and 128.1 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 29.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations and maintenance. Feed to the TDF was stopped at 0926 hours on Tuesday, March 16, 1999 in order to replace a sprocket inside of the Wet Ash Drag (WAD). Hazardous waste feed was resumed at 1456 hours once the repairs were complete. The contractor also performed Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris handling, wood debris management, equipment decontamination, lime mixing, shot rock backfilling, decontaminated rock backfilling, ash backfilling and ash handling. Excavation activities were continued directly south of the WWTP, in the vicinity of former sediment basin 2. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: At 1653 hours on Tuesday, March 16, 1999, waste feed to the TDF was stopped because of a problem with one of the kiln's trunion bearings which caused the kiln to temporarily stop rotating. Hazardous waste feed was shut off and a redundant drive motor started in order to keep the kiln rotating. The contractor then retracted the dual feed screws in order to allow atmospheric air to cool the system. While performing the required lockout/tagout of the feed screws, electrical power to the Induced Draft (ID) fan was turned off, initiating a Thermal Relief Valve (TRV) opening at 1737 hours. Feed soil had been suspended for 44 minutes prior to the TRV opening which lasted for approximately 60 seconds. Hazardous waste feed to the incinerator will remain off-line until the trunion bearing can be repaired and re-installed. At this point work is expected to last about 4-5 days and waste feed to the TDF is anticipated to resume early next week. The fourth and final stack testing event for metals and particulate will be rescheduled for a later date.

V. NOTED VISITORS: Tony Eicher (FOCUS).

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Tuesday, March 16, 1999

FOR DATES: From 7:00 a.m. Monday, March 15, 1999 to 7:00 a.m. Tuesday, March 16, 1999.

WEATHER: Partly cloudy. The temperature range was 26oF to 38oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 705.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 29.7 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris handling, wood debris management, lime mixing and ash handling. Excavation activities were continued directly south of the WWTP, in the vicinity of former sediment basin 2. As the excavation progressed across the face of the sheet piling, shot rock was backfilled directly behind. Continued with the real time perimeter air monitoring (PAM) program and collected the OSHA method 93 air sampling media at the on-site and off-site air monitoring stations.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. The fourth and final quarterly stack testing event for metals and particulates is scheduled for March 17th and 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Tony Eicher (FOCUS).

PREPARED BY: Steve Gibson for William Werntges (USACE)


Monday, March 15, 1999

FOR DATES: From 7:00 a.m. Sunday, March 14, 1999 to 7:00 a.m. Monday, March 15, 1999.

WEATHER: Overcast, snow starting about 1800 hrs. The temperature range was 27oF to 36oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 725 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 30.3 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris handling, wood debris management, lime mixing, ash handling and ash backfilling. Excavation continued in the northeast corner of the exclusion zone, where the sheet pile along the R&R tracks intersects the sheet pile which runs along the southern end of the WWTP. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-3-6-99-380) for the contents of bin 1. The analytical report indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. The fourth and final quarterly stack testing event for metals and particulates is scheduled for March 17th and 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for Chad Thompson (USACE)


Sunday, March 14, 1999

FOR DATES: From 7:00 a.m. Saturday, March 13, 1999 to 7:00 a.m. Sunday, March 14, 1999.

WEATHER: Partly cloudy. The temperature range was 23oF to 44oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The incinerator remained in a hot hold until maintenance activities were completed at 2034 hrs. on 3-13-99, at which time the processing of site soils was resumed. The TDF was on-line for 10.4 hours and 271.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 26.1tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations and maintenance. Maintenance continued on the ash drag conveyors and the steam scrubber duct work . The contractor also continued Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, equipment decontamination, debris handling, debris separation, wood debris management, lime mixing, ash handling and ash backfilling. Excavation was performed along the railroad sheet-piling in the vicinity of former sediment basin 2, and minor excavation was performed inside of the feed preparation building. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. The fourth and final quarterly stack testing event for metals and particulates is scheduled for March 17th and 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Richard Lupold for William Werntges (USACE)


Saturday, March 13, 1999

FOR DATES: From 7:00 a.m. Friday, March 12, 1999 to 7:00 a.m. Saturday, March 13, 1999.

WEATHER: Partly cloudy. The temperature range was 24oF to 39oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The incinerator is currently in hot hold while maintenance is continued on the bottom ash drag conveyors. The TDF was on-line for 1.0 hours and 28.6 tons of contaminated soils were thermally treated during the time frame of this report. The unit was taken off-line at 0800 hours on Friday, March 12, 1999 in order to perform the monthly interlock checks, the monthly instrument calibrations and to perform maintenance on TDF equipment. The TDF processed contaminated soil at an average rate of 28.6 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations and maintenance. TDF maintenance activities included the replacement of several flights on both ash drags and several sprockets on the Wet Ash Drag (WAD), the repair of some internal water leaks on the Dry Ash Drag (DAD) floor plates, and the cleaning of the steam scrubber's duct work. The contractor also continued Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, equipment decontamination, debris handling, debris separation, wood debris management, lime mixing, rock cleaning, rock backfilling, ash handling and ash backfilling. Excavation was performed along the railroad sheetpiling in the vicinity of former sediment basin 2, and minor excavation was performed inside of the feed preparation building. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. The fourth and final quarterly stack testing event for metals and particulates is scheduled for March 17th and 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Friday, March 12, 1999

FOR DATES: From 7:00 a.m. Thursday, March 11, 1999 to 7:00 a.m. Friday, March 12, 1999.

WEATHER: Sunny and clear. The temperature range was 20oF to 38oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 20.0 hours and 491.7 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is attributed to re-heating the TDF to operating temperatures after a brief power outage, which caused the waste feed and all of the burners to shut off. The TDF processed contaminated soil at an average rate of 24.6 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

On this date, feed to the incinerator was stopped at approximately 0800 hours. The TDF is currently in hot hold in order to perform the monthly interlock checks (AWFSOs) and perform the monthly calibrations to the TDF monitoring equipment. In addition to the AWFSOs and the monthly calibrations, some preventative maintenance will be performed while the unit is off-line. The TDF is planned to be off-line for approximately 8-hours.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris handling, wood debris management, lime mixing, ash handling and ash backfilling. Excavation activities were performed in the vicinity of sediment basin #2 and inside of the feed preparation building. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical reports (bot-ash-3-3-99-378A and bot-ash-3-3-99-378C) for the contents of bin 3, and analytical report (bot-ash-3-5-99-379) for the contents of bin 2. Each of the analytical reports indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site air monitoring, perimeter air monitoring, WWTP operations, rock cleaning, inorganic debris decontamination and the implementation of the site close out plan. The fourth and final quarterly stack testing event for metals and particulates is scheduled for March 17th, and 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: A defect in a new motor used on a plant air compressor motor caused a momentary power interruption on the TDF pad, resulting in an unplanned TRV opening at 1240 hours on Thursday, March 11, 1999. Power to the unit was restored and feed resumed at 1644 hours, once the unit reached operating temperatures. The proper local authorities were notified of the TRV opening.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Thursday, March 11, 1999

FOR DATES: From 7:00 a.m. Wednesday, March 10, 1999 to 7:00 a.m. Thursday, March 11, 1999.

WEATHER: Partly sunny, cold. The temperature range was 20oF to 39oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 554.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 23.3 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris handling, wood debris management, lime mixing, sheetpiling monitoring and ash handling. Imported shot rock was placed as part of the 2-foot rock bottom. Excavation activities were performed in the vicinity of sediment basin #2. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site air monitoring, perimeter air monitoring, WWTP operations, rock cleaning, inorganic debris decontamination and the implementation of the site close out plan. The fourth and final quarterly stack testing event for metals and particulate is scheduled for March 17th, and 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Chad Thompson for Frank Cleaver (USACE)


Wednesday, March 10, 1999

FOR DATES: From 7:00 a.m. Tuesday, March 9, 1999 to 7:00 a.m. Wednesday, March 10, 1999.

WEATHER: Overcast with snow. The temperature range was 18oF to 25oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 566.3 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 23.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris handling, wood debris management, lime mixing, sheetpiling monitoring and ash handling. Imported shot rock was placed as part of the 2-foot rock bottom. Excavation activities were performed in the vicinity of sediment basin #2. Continued with the real time perimeter air monitoring (PAM) program and collected the weekly OSHA method 93 air sampling media at the on-site and off-site air monitoring stations.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site air monitoring, perimeter air monitoring, WWTP operations, rock cleaning, inorganic debris decontamination and the implementation of the site close out plan. The fourth and final quarterly stack testing event for metals and particulate is scheduled for March 17th, and 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Mike Welch and Larry Newcomer (PADEP).

PREPARED BY: Steve Gibson for William Werntges (USACE)


Tuesday, March 9, 1999

FOR DATES: From 7:00 a.m. Monday, March 8, 1999 to 7:00 a.m. Tuesday, March 9, 1999.

WEATHER: Clear and sunny. The temperature range was 6oF to 33oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.6 hours and 519.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 22.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris handling, lime mixing and ash handling. Excavation activities were performed in the vicinity of sediment basin #2, the sediment basin is being eliminated as part of the exclusion zone close out. Continued with the real time perimeter air monitoring (PAM) program and began the weekly OSHA method 93 24-hour air sampling event for BNA at the on-site and off-site air monitoring stations.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Monday, March 8, 1999

FOR DATES: From 7:00 a.m. Sunday, March 7, 1999 to 7:00 a.m. Monday, March 8, 1999. WEATHER: Overcast. The temperature range was 6oF to 28oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 522.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 21.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, excavation, feed preparation, debris handling, lime mixing, ash handling and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for Chad Thompson (USACE)


Sunday, March 7, 1999

FOR DATES: From 7:00 a.m. Saturday, March 6, 1999 to 7:00 a.m. Sunday, March 7, 1999.

WEATHER: Heavy snow and rain. The temperature range was 16oF to 35oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 495.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 20.7 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, equipment decontamination, debris handling, debris separation, wood debris management, lime mixing, ash handling and ash backfilling. Minor excavation was performed in the center of the site. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical reports (bot-ash-02-26-99-375B and bot-ash-02-26-99-375A) for the contents of bin 6. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Saturday, March 6, 1999

FOR DATES: From 7:00 a.m. Friday, March 5, 1999 to 7:00 a.m. Saturday, March 6, 1999.

WEATHER: Partly sunny. The temperature range was 17oF to 34oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 496.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 20.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, equipment decontamination, debris handling, debris separation, wood debris management, lime mixing, ash handling and ash backfilling. Minor excavation was performed along the containment berm on the southern end of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Mike Welch and Sandra Downs (PADEP).

PREPARED BY: Steve Gibson for William Werntges (USACE)


Friday, March 5, 1999

FOR DATES: From 7:00 a.m. Thursday, March 4, 1999 to 7:00 a.m. Friday, March 5, 1999.

WEATHER: Overcast with snow, clear and windy in the afternoon. The temperature range was 14oF to 33oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line 17.7 hours and 373.8 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is attributed to raising the unit to operating temperature after the two TRV openings experienced during the morning hours. Once the kiln reached operating temperature, the unit was placed in 'hot hold' while maintenance was performed on the double dump valve. The TDF processed contaminated soil at an average rate of 21.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations and maintenance, hazardous waste feed to the incinerator was brought back on-line at 1318 hours on Thursday, March 4, 1999. The contractor continued Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris handling, debris separation, lime mixing, wood debris management, ash backfilling and ash handling. Continued with the real time perimeter air monitoring (PAM) program. Performed the weekly NPDES water-sampling event in the WWTP. No new excavation was performed on this date. One roll-off containing PPE and decontaminated miscellaneous metal was shipped off-site as non-hazardous waste to a hazardous waste landfill.

The TDF experienced two brief Thermal Relief Valve (TRV) openings during the morning hours of Thursday, March 4, 1999, one at 0725 hours for 80 seconds, and one 0743 for 89 seconds. Both of the TRV openings occurred during two separate local power fluctuations experienced on-site, evidently caused by the weather conditions Thursday morning. Prior to the TRV opening, feed had been off-line for approximately 12 hours. The proper local authorities were notified of the TRV opening.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and implementation of site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Thursday, March 4, 1999

FOR DATES: From 7:00 a.m. Wednesday, March 3, 1999 to 7:00 a.m. Thursday, March 4, 1999.

WEATHER: Overcast with heavy, approximately 2", rain, later changing to snow accumulating approximately 4" by the close of this report. The temperature range was 29oF to 57oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 11.9 hours and 264.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 22.3 tons per hour while on line. Off-line time is attributed to the intentional stoppage of feed due to electrical storms and burner system maintenance. Feed was stopped at 7:13 P.M. last evening per site SOP as a precaution to a potential loss of power due to the electrical storm. During this down time, maintenance to a thermocouple in one of the kiln burners was begun and feed has not resumed as of 7:00 A.M. today. Related ongoing site activities include snow removal, feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris handling, debris separation, lime mixing, wood debris management, ash backfilling and ash handling. Continued with the real time perimeter air monitoring (PAM) program and finished the weekly OSHA method 93 tests for BNA at the on-site and off-site air monitoring stations. Excavation activities were continued directly south of sediment basin #2.

Received bottom ash analytical report (bot-ash-02-23-99-373) for the contents of bin 1 and received reports (bot-ash-02-24-99-374A and bot-ash-02-24-99-374B) for the contents of bin 7. Ash analytical reports indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site. Received fly ash analytical report (fly-ash-02-23-99-074) for the contents of bin B. The analytical laboratory test results indicated that the concentration of chromium was greater than 25 times the PADEP drinking water standard. Therefore the contents of bin B will be stabilized, re-sampled and re-analyzed prior to backfilling.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and implementation of site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: The TDF experienced two brief Thermal Relief Valve (TRV) openings during the morning hours of Thursday, March 4, 1999, one at 7:25 A.M. for 80 seconds, and one at 7:42 A.M. for 89 seconds. Both of the TRV openings occurred during two separate local power fluctuations experienced on-site, evidently caused by the current weather conditions. Prior to the TRV opening, feed had been off-line for approximately 12 hours. The proper local authorities were notified of the TRV opening.

V. NOTED VISITORS: Pearl Lamey, representing the local environmental advisory committee, was given a tour of the site.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Wednesday, March 3, 1999

FOR DATES: From 7:00 a.m. Tuesday, March 2, 1999 to 7:00 a.m. Wednesday, March 3, 1999.

WEATHER: Clear & sunny. The temperature range was 32oF to 46oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 540.1 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 22.6 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris handling, debris separation, lime mixing, wood debris management and ash handling. Excavation activities were continued south of sedimentation basin 2. Eliminated the southern 1/3 of sedimentation basin 2. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and implementation of site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Larry Newcomer, PADEP

PREPARED BY: Danny Zion for Chad Thompson (USACE)


Tuesday, March 2, 1999

FOR DATES: From 7:00 a.m. Monday, March 1, 1999 to 7:00 a.m. Tuesday, March 2, 1999.

WEATHER: Partly cloudy. The temperature range was 31oF to 39oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.6 hours and 542.8 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is primarily attributed to the automatic stoppage of feed during a flameout of one of the Secondary Combustion Chamber (SCC) burners. The TDF processed contaminated soil at an average rate of 23.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris handling, debris separation, lime mixing, wood debris management and ash handling. Excavation activities were continued on the eastern side of the exclusion zone. The contractor continued to implement the wastewater management close out plan. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and implementation of site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Monday, March 1, 1999

FOR DATES: From 7:00 a.m. Sunday, Feb. 28, 1999 to 7:00 a.m. Monday, March 1, 1999.

WEATHER: Overcast, rain. The temperature range was 37oF to 40oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 566.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 23.7 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris handling, debris separation, lime mixing, wood debris management, 2' bottom rock placement, and ash handling. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and implementation of site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for Chad Thompson (USACE)

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.